OpenStudy (anonymous):

.a square pyramid with base edge 80 m and slant height 100 m 44,800 m2 48,000 m2 16,000 m2 22,400 m2

jimthompson5910 (jim_thompson5910):

Area of base: A = s^2 A = (80)^2 A = 6400 --------------------- Area of one triangular face: A = (bh)/2 A = (80*100)/2 A = 8000/2 A = 4000 So the area of the four lateral triangular faces is 4*4000 = 16000 Add this to the area of the base to get 16000 + 6400 = 22400 So the total surface area of this pyramid is 22400 square meters.

OpenStudy (anonymous):

Thanks :)

jimthompson5910 (jim_thompson5910):

yw